OpenStudy (steve816):

Please help me! Find x so that the distance between the points (6, -1) and (x, 9) is 12.

OpenStudy (steve816):

So far, this is what I have:\[12=\sqrt{x^2-12x+138}\] Not sure what to do from here :/

ganeshie8 (ganeshie8):

still here ?

OpenStudy (steve816):

Yes I'm here!

jimthompson5910 (jim_thompson5910):

\[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] \[\Large 12 = \sqrt{\left(x-6\right)^2+\left(-1-9\right)^2}\] \[\Large 12 = \sqrt{\left(x-6\right)^2+\left(-10\right)^2}\] \[\Large 12 = \sqrt{x^2-12x+36+100}\] \[\Large 12 = \sqrt{x^2-12x+136}\] I'm not sure how you got 138. Did you mean 136??

OpenStudy (steve816):

Yeah, I meant 136

OpenStudy (steve816):

@ganeshie8 Please! I need a step by step guide.

ganeshie8 (ganeshie8):

square both sides, what do you get ?

ganeshie8 (ganeshie8):

\(12 = \sqrt{x^2-12x+136}\) square both sides and get \(144 = x^2-12x+136\) rearranging a bit \(x^2-12x-8=0\)

ganeshie8 (ganeshie8):

you may use quadratic formula to solve x

OpenStudy (steve816):

Thank you ganeshie!